UT Feb 20 2022 - Split Night PEPSI/PETS and DEC



Staff

  • OSA: S.Allanson
  • ISp: Mtn Staff
  • Mtn Manager: R.Hansen
  • ISA/Observer: B.Rothberg
  • AO: J.Power
  • Special Guests: R.T. Gatoo

Night Info

  • Sunset 18:09
  • Nautical Twilight Ends 18:34
  • Astro Twilight Ends 19:31
  • Moon rise 21:16
  • Handover from LBTO/DEC to PETS/PEPSI 22:20
  • Handover from PETS/PEPSI to LBTO/DEC 03:11
  • Astro Twilight Starts 05:35
  • Nautical Twilight Starts 06:04
  • Sunrise 06:57
  • Moon 89.5%

Night Plan

  • Pending any updates from AO SW and AO group - install test ao supervisor update
  • Starting Config Bino-LUCIs - Authorize LUCI-1
  • AO/ESM Scripts
    • /home/lbto/Service_Observing/2022A_D_E/LUCI/AO_ESM/ESM_V36.xml
    • /home/lbto/Service_Observing/2022A_D_E/LUCI/AO_ESM/AO_Testing_V73.xml
  • DONE OSCO until 19:31
  • DONE LUCI-1 AO Science - 1hr - Zhang
    • /home/lbto/Service_Observing/2022A_D_E/LUCI/XZhang_LD2022A_0001/Zhang_LD2022A-0001_V3.xml
    • On-axis R=14.6 (15.8) Bin 3
    • Seeing <1.2"
    • thin clouds
    • Observe science target first - if clearly resolved complete
    • If unresolved or extended - observe PSF star
  • Check Bin 1 Full AO IQ - at ~360 deg instrument rotator position - 0.75hr
    • Calculate and select possible targets during night - want something around 360
    • Quick: Rotangle = PA - PARANGLE + 7.735
  • Full AO test w/N375 Camera - 1hr
    • characterize degradation in corrections as function of distance from AO Ref star
    • Bin 1/2 and Forced Bin 4
      • M34 - twilight-22hr
      • DONE M67 - 19:30-03:15hr
      • DONE M92 - 02:20-twilight
  • DONE PEPSI/PETS 22:20-03:11hr
  • Moon 15deg Bin 4 test -0.75hr
    • DONE RA12hr-13hr folder Star#9 R=16.7(17.9) HS 1338+0807 WD
  • AO "Pretty Pix"
    • Cat's Eye Nebula - On axis
    • R=11.38 (12.58 - Bin 1 or Bin 2 - borderlien depending on WFSmag)
    • Visible 03hr-twilight
    • DONE H2, semi-DONE FeII, P-Beta, Ks
  • DONE Roll back ao-supervisor "latest" version after closing up
  • Other AO tasks
    • Bin 1 ESM - M61 R=10.47 - face on spiral
    • visible handover-twilight

Backup

  • Poor Weather
    • First half backups
      • CaT RQQ Galaxies
      • HS 0624+6907 - bino LUCI
      • Same condition requirements as mergers - visible twilight-3hr
    • 2nd half backups
      • Local Mergers - Rothberg
        • /home/lbto/Service_Observing/2022A_D_E/PoorWeather/Rothberg/LocalMergers/COBandhead_Mergers_V41.xml
        • Any conditions (guiders up to 4" - as long as guider can lock on and WFS) - no particular order
          • UGC 9829 -2nd half
          • UGC 10607 -2nd half
          • UGC 10675 -2nd half
          • NGC 4004 -2nd half
    • If clear and super poor-poor seeing (regardless of time of night)
      • LUCI-Bino Standards
        • LUCI - /home/lbto/Service_Observing/2022A_D_E/PoorWeather/LUCI_Phot_Standards/LUCI_Phot_ZP_OT_V3.xml
        • If seeing > 4" - make these track modes in the OT and export
  • If AO cannot be executed and conditions are good
    • LUCI-Bino SL - GRB 211211A - 1.25hr (2nd half) - Rothberg
      • OT File: /home/lbto/Service_Observing/2022A_D_E/LUCI/GRB211211A/GRB211211A_LD2021B-002_V5_OT.xml
      • Ks and J (fraternal) imaging
      • Seeing < 1.3"
      • Thin Clouds
      • If clear (photometric) - execute standard GSPC P272-D before and after observations (just rotates through Ks and J)

Night Log

Closed Dome

  • UT 00:15 - initializing LUCIs, taking test darks
  • L1 0001-0002; L2 0001-0002 -2.51sec dark
  • L1 0003-0007 check N30fieldstop
  • UT 00:25 - LUCIs ready

On-Sky

  • UT01:08 - opening dome
  • UT 01:24 - pointing preset
  • UT 01:33 - collimation preset
  • UT 01:35 - SX guider shows 0.7"
OSCO
  • UT 01:39 sending preset to OSCO star HD 38515 (R=7.49, WFSmag = 8.8)
  • GS too close to the edge, choosing new guidestar, regenerating script - IT 8495 for AGW patrol field
  • UT 01:43 - acquiring
    • L1 0008-0017 NCPA off Instument rotator = 336
    • L1 0018-0027 NCPA off instrument rotator = 335.7 - 2nd dither position
    • No NCPA PSF looks really good
    • Bin1NoNCPA.png
  • Reversing offset UT 01:56
    • NCPA on
    • L1 0028-0036 Instrument rotator = 334
    • Bin1wNCPA2.png
    • Not sure if much difference, but it does not look signififcantly worse with NCPA on - there is a darker spot in the first ring at 11 o'clock
    • L1 0037-0046 offset intrumenet rotator = 332
    • OSCO completed

Science Zhang

  • UT 02:06 - sending preset
  • Instrument rotator = 214
  • Rmag = 14.64, WFSmag = 14.4 - shifting to Bin 2
  • UT 02:11 - is the WFS not updating? - probe position may have drifted error - cancelled AO perset - IT 8210, IT 8495
  • UT 02:13 - resending preset - probe position OK now
  • CCD47 star looks elongated
  • UT 02:16 - CCD47 also shows a 2nd source 6" away
  • UT 02:17 - Optimize gain failed first time, retrying
  • Safeskipping - HO2 ramping up to 15
  • NCPA is on
    • L1 0047-0048 sub-window acquisition frames 1012,1085 center of frame
  • UT 02:24 - starting script
  • distance between objects is 0.59" - relative mag difference between the two objects ~1.5mag
    • TargetZhang.png
  • UT 02:26 - safeskipping - paused - J.Power reclosing loops, setting gains manually (took us 10min to close loop and auto-optimize)
    • L1 0049 safeskipping - cube may be impacted - will take dither position again
    • as soon as HO2 is applied, safeskipping occurs and a weird shape on mirror - camera lens is not moving (we are in Bin2)
    • GOpt may be the problem - may not have exited gracefully - when loops opened it didn't turn off - couldn't turn off as another rogue GOpt window was open
    • only way to reset is to optimize or send a new preset - not opening the loops
  • UT 02:42 - closed loop and started (gains are lower)
  • some elongation in some frames
    • L1 0050-0052 - Ks
  • 2nd dither too on edge #0052
  • Redoing script with +/-0.75" DETX offsets
  • Restarting from 2nd dither on Ks
    • L1 0053-0054 Ks (with new dither)
  • first H-band looks poorer IQ
  • UT 02:56 - paused before last H-band image - increased tiptilt
    • L1 0055-0057 - H
    • L1 0056- H-band (looks a little poorer IQ)
  • UT 03:00 - pausing script - re-optimizing gain
    • L1 0058 J-band
  • UT 03:08 - redid gains, closed loop re-doing H-band
    • L1 0059-0062 H-band
    • L1 0061- loops opened - looks like pupils were misaligned
  • UT 03:18 - loops now re-closing
    • L1 0062 H-band
    • L1 0063 J-band
    • L1 0064 - J-band but loops opened - BAD IMAGE
  • UT 03:23 - loops opened again - re-opimizing
    • Closing GEIRS session - became useless for real time monitoring of PSF
  • UT 03:29 - loops closed
  • UT 03:31 - loops paused - offset got stuck in IIF - fortunately when I relaunched GEIRS it undid sub-frame so aborting that and redoing took time
  • UT 03:32 - resuming script from 2nd J-band dither
  • UT 03:33 - ignore LUCI images 0065-0074 - GEIRS reset the subframe
  • UT 03:34 -
    • L1 0075-0076 J-band
    • L1 0077 Y1-band
  • UT 03:38 - Safeskipping, loops opened, LUCI IIF offset command failed
  • UT 03:40 - re-optimizing
  • UT 03:44 - resuming script 2nd dither Y1
    • L1 0078-0079 Y1-band
  • UT 03:47 - tip-tilt off-loading not enabled
    • offsets are taking longer
  • UT 03:48 - offset completed
    • L1 0080-0082 - Y2-band
PSF Star for Zhang Science UCAC4 704-031301
  • UT 03:58 - sending preset for PSF Star Rmag=14.38, WFSmag = 14.5 NCPA On
  • Subframe center 1012,1110
  • L1 0083-0084 - subframe acquisition
  • UT 04:09 - staring exposures
    • L1 0085-0087 - Ks
    • L1 0088-0090 - H-band
    • L1 0091-0093 - J-band
    • L1 0094-0096 - Y1-band
    • L1 0097-0099 - Y2-band
  • UT 04:36 complete

N375 Camera with Full AO - M67

  • UT 04:37 - sending preset, Rmag=10.73, WFSMag = 12.3
  • UT 04:39 - aborting script, J.Power resetting gain ramp to 0ms (as originally set) - restarting arbitrator
  • UT 04:46 - canclled LUCI preset - have noise on OCAM - need to rebin in Bin2 then rebin again in Bin 1
  • UT 04:51 - re-sending preset - changed K to BrGamma to avoid saturation (did this while J.Power was working on noisy OCAM issue)
  • UT 04:54 - pupils were railing, finally caught, instrument rotator = 224
  • UT 04:58 - DIMM shows 0.9"
  • UT 05:00 - safeskipping
  • UT 05:01 - NCPA on, only way to get rid of noise is power-cycle OCAM which would take 20min
    • L1 0100-0107
    • Aborted image exposure is going but loops are paused - 100%safeskip
  • UT 05:21 - putting LUCI-1 in safe config - ready for PEPSI reconfig

PEPSI/PETS-6

  • UT 07:48 - wind hitting limits at the SX ASM - approaching 6 m/s - almost had to abort observation
  • UT 08:01 - keeping an eye on winds -still facing directly into the wind
  • UT 08:08 - moving to telescope off source - we have reached the 6 m/s limit
  • UT 08:11 - moving off TOI 561 to another target rho Virgo - will stay about 30min - then move back to transit to check winds
  • UT 08:54 - moving back to TOI 561 -- worked
  • UT 10:11 - PETS complete

LUCI-1 AO/ESM Testing Continues

  • UT 10:11 reconfiguring to mono-LUCI
  • UT 10:13 - moving N30Fieldstop into place
  • UT 10:16 - checking Fieldstop L1 0108-0110
  • UT 10:20 - pointing check

Moon Test -15deg away Bin 4 HS 1338+0807

  • UT 10:22 -sending preset, Rmag=16.7, WFSmag= WD star
  • UT 10:24 - no WFS on AGW1 yet... - seems to be the probe drift error - cancelled AO preset, aborted LUCI script - IT 8210
  • UT 10:26 - resending preset
  • UT 10:28 - OCAM looks noisey again (not counts from Moon, electronics noise same as before - LL and UR quadrants bad) - centered pupil centered on noise and drove things off the hotspot
  • UT 10:30 - taking new dark for OCAM
  • UT 10:31 - changing binnings on OCAM (Bin 3 and back to Bin 4) and back to try and get rid of noise
  • UT 10:33 - cycling OCAM, the bin swap didn't fix it -IT 8288 (now have seen this in Bin 1 and Bin 4 and NOT at the start of the night)
  • UT 10:36 - Confirming fieldstop alignment while OCAM gets its act together -
    • Noticed several times now, if RESET ALL and STOP on RTD wihile on AFC tab it says "LMS file not found" - reset and STOP START because the pixel histogram button wouldn't launch the pop-up strech window
    • L1 0111-0113
  • UT 10:44 - noise is gone
  • UT 10:51 - sending preset
  • UT 10:55 - star is wandering off - but center star got it back, now centering pupils
  • UT 10:56 - J.Power manually adjusting
  • UT 10:57 - resending preset -
    • reason: center pupil wandered off, which pulled the guidestar off the hostpost far enough that SH spots were empty - if we had closed the loop, then the "feature" where when the off-axis WFS is paused would have paused AO WFS
  • UT 10:58 - pupils still way off -almost an entire pupil distance over
    • can't move the camera lens manually enough to catch the pupils - is this the wrong pupil for Bin4?
    • now it grabbed it - had to rail camera lens to catch the pupils
  • UT 11:09 - pretty terrible PSF - FWHM~0.3" (target is H=17.1 according to 2MASS) - DIMM 0.7"
  • Hband_image_Bin4_15degMoon_H17pt1.png
  • PSF of star is 0".24-0".3
  • on the offset we see this in IIF
  • Sun Feb 20 11:15:11.736 2022 left Offset mount delay for LUCIFER: 15132 milliseconds (minimum 0)
  • Fell back to active mode - but I set it up in OT at this PA not to - looking at OT again -its just on the edge
    • safeskipping loop opened during frame #0121
    • pair subtracted image FWHM 14pix x 12 pix (from RTD) 0".21 x 0".18
  • Last offset - lots of safeskipping (#0124)
    • when we got back from 100" offset it started doing active optics not the AO WFS - why did this happen? we went back on-source and the star was on the OCAM

M92 Full N375 Test

  • UT 11:26 - sending preset
  • UT 11:30 - star in CCD47 is super faint - cancelling preset - doing pointing check
    • we've used this star before for ESM Bin in same field, so likely pointing error
  • UT 11:32 - resending preset - weren't far off on pointing check
  • Rmag=9.43 (UCAC4 says 12.0)
  • UT 11:36 - sending new version with AO Ref star nearby R=11.73 - might be a companion 4" away (not shocking this is a Globular Cluster)
    • in Bin2
  • UT 11;38 - OCAM noise in "sectors" appeared again
    • Noise vanished when it auto-switched to Bin2
  • we might be in a seeing bubble
  • UT 11:46 - DIMM back down to <1"
  • Offsets back from sky are taking a while
  • UT 11:53 - PROBE Failed to move to offset - now coming back
  • Sun Feb 20 11:53:02.727 2022 left Offset mount delay for LUCIFER: 131754 milliseconds (minimum 0)
  • Reverted back to active preset after error above - no AO ref star came back into CCD 47 - did the bayside stages not move correctly? - they did try - the command was to a position beyond the patrol field
  • According to OT we were NOT at the edge, but near it - changed this by 10" DETX in V3 of script and it worked - need to update OT with changed AO Patrol FIeld
  • Getting FWHMs 2.3-2.8pix full across the FOV - is this SUPER-ESM?
    • L10134 -safe skipping - this is after a large offset - do not use this image
    • L1 0125-0134

Cat's Eye Nebula

  • UT 12:16 - sending preset
  • Guiders shows 0.8" before closing loop
  • Rmag = 11.38, WFSman = 13.0
  • UT 12:23 - another seeing bubble - jumped to 1.3" on DIMM
  • UT 12:24 - started -
  • UT 12:26 - fixing N30 fieldstop (had a pause)
    • L1 0135 - 2.51 sec central pixel of star is 52k in Ks
  • Took a while to align L1 0135-0138 w/Ks
  • UT 12:29 - staring real part of script with H2
  • 1500 counts 3sec H2 - changing script to 30sec for H2 - central pixel 20k
    • FWHM~0.056"
  • UT 12:36 - shell failure - failed before the return from skyframe
    • CCD47 went not-live
  • A resumeAO was sent when we were on-sky
    • L1 0139 - H2 3DIT x20NDIT
    • L1 0140-0143 - H2 30DIT x 2NDIT
    • Below is 30sec x 2 collapsed image from the RTD H2 Cats Eye
    • Catseye_H2_1frame.png
  • UT 12:43 - J.Power recovering ASM falillure
  • UT 12:45 - resending preset - set Rmag =13.0 to save time
    • guider shows FWHM=0.71"
    • L1 0144- 2.51sec Ks test image
    • L1 0145-H2 (1st exposure repeats 0,0 position)
    • L1 0146-0151 - H2
  • UT 13:05 - OCAM not live on offset - ASM failure for off
  • UT 13:10 - DIMM ~1"
  • UT 13:12 - sending preset again - AGW shows 0.67" before closing loop
  • UT 13:18 - closed loop
    • L1 0152 - Ks 2.51sec
    • L1 0153-0155 H2
    • L1 0156-0159 FeII (0.048")
  • UT 13:32 - safeskip - guiding on sky - aborted exposure
  • UT 13:32 - putting LUCI-1 in safe config - closing up
AO supervisor was reverted by J.Power

UT 14:45 - putting LUCI-1 into safe config


PROBLEMS

AO

  • Bin 2 Full AO appears to be unstable now - can't keep loops closed - they open every few minutes - need to keep reoptimizing gains
    • Target was R=14.6, closed as R=14.4, 3.4 photons/subapp
    • seeing is steady DIMM reading 0.7", offa-axis guiders showing 0.5"-0.6"
    • could the AO supervisor test be causing all these problems?
    • AO operator unable to change the gains, gain vector was empty during above events - not clear if this is a bug from update or (as normal) gain vectors are empty during 100% safeskip or offset
    • JC Guerra suggested during software ao supervisor version swap changing the offset to 2000ms from 0ms - feels like this causing longer offsets or porblems
      • J.Power speculated GOpt GUI the 2000ms addition is causing 2seconds (or more) longer on the offsets
      • Set back to 0ms - loop returned simultaneously with gain vector loading
    • Could AO supervisor change be triggering more open loops?
    • PSF star is now more stable - no issues - companion star was 5x fainter and 0.6" away - X.Zhang suggested this could have added instabilities?
  • Offsets - seem to be taking longer - getting several timeouts on IIF when coming back from sky offsets - this is after setting the loop return to 0ms
    • could also be influenced by the AGW guideprobe problems (see ITs)
  • During UT 12:36 failure a "resumeAO' command was issued (no one did it) while we were on the sky frame
    • PauseAO was never successfully completed
  • UT 13:05 - OCAM not live - ASM failure - not sure which triggered the other - needs to be investigated
    • seemed to be triggered by large offset to sky - but need to verify
    • confusion in the current state - so OCAM may have gone not live


LUCI-1 Cals

  • UT 13:43 - starting N30 H2 imaging flats
    • L1 0160-0169 - fieldstop shifted again - need to adjust - DO NOT USE
    • L1 0170 -0171 adjusting Fieldstop Ks
    • L1 0172-0181
  • UT 13:56- starting N30 FeII imaging flats
    • L1 0182-0191
  • UT 14:01 - starting N30 Ks imaging flats
    • L1 0192-0201
  • UT 14:08 - starting N30 H imaging flats
    • L1 0202-0211
  • UT 14:14 - starting N30 J imaging flats
    • L1 0212-0221
  • UT 14:22 - staring N30 Y1 Imaging flats
    • L1 0222-0231
  • UT 14:26 - starting N30 Y2 imaging flats
    • L1 0232-0241
  • UT 14:30 - starting N375 BrGamma Imaging Flats
    • L1 0242-0251
  • UT 14:36 - starting N375 K Imaging Flats
    • L1 0252-0261


-- BarryRothberg - 19 Feb 2022
I Attachment Action Size Date Who Comment
Bin1NoNCPA.pngpng Bin1NoNCPA.png manage 28 K 20 Feb 2022 - 01:55 BarryRothberg  
Bin1wNCPA.pngpng Bin1wNCPA.png manage 69 K 20 Feb 2022 - 02:00 BarryRothberg  
Bin1wNCPA2.pngpng Bin1wNCPA2.png manage 34 K 20 Feb 2022 - 02:02 BarryRothberg  
Catseye_H2_1frame.pngpng Catseye_H2_1frame.png manage 337 K 20 Feb 2022 - 12:42 BarryRothberg  
Hband_image_Bin4_15degMoon_H17pt1.pngpng Hband_image_Bin4_15degMoon_H17pt1.png manage 899 K 20 Feb 2022 - 11:12 BarryRothberg  
TargetZhang.pngpng TargetZhang.png manage 74 K 20 Feb 2022 - 02:27 BarryRothberg  
Topic revision: r4 - 01 Mar 2022, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback